# A Comprehensive Benchmark for RNA 3D Structure-Function Modeling

### v1

@@ 6. RNA-Ligand Pocket Multi-class (3) 203-43-44
4.1 RNA-GO: Function Tagging
(Definition): This is a multi-label classification task where input RNA is mapped to a sequence of 5 possible labels representing molecular functions.
(Context): Effective function prediction models have the capacity to uncover new structure-function connections. The Gene Ontology (GO) (Ashburner et al., 2000) was developed to associate a function to a gene, and thus, indirectly to the RNA or protein it encodes. It resulted in discrete functional categories called GO terms. Predicting the GO term from a protein structure was proposed as a task in (Gligorijević et al., 2021) and has since been regularly used as a benchmark task. A mapping between RNA sequences and GO terms is available in Rfam (Griffiths-Jones et al., 2003; Ontiveros-Palacios et al., 2025), a database of non-coding RNA families which are manually annotated with GO terms.
(Processing): We extracted all GO terms corresponding to an RNA structure. Then, we filtered these terms based on their frequency, removing the two most frequent labels (ribosomal systems and tRNAs), as well as infrequent ones (less than 50 occurrences). We then grouped together GO-terms that were fully correlated, which resulted in five classes. Finally, we discarded RNA fragments with less than 15 residues, resulting in 501 systems. We split these systems based on a sequence similarity cutoff, following Gligorijević et al. (2021). Accuracy, AuROC and the Jaccard index are the metrics computed for evaluation.
4.2 RNA-IF: Molecular Design
(Definition): This is a residue-level, classification task where given an RNA structure, we mask all residue identities (by keeping only the coordinates of the backbone) and aim to learn the mapping from structure to sequence.
(Context): One avenue for designing new macromolecules is to design a coarse grained tertiary structure first, for instance scaffolding a structural motif, and then design a sequence that will fold into this given structure. This second step is denoted as Inverse Folding (IF), since it maps a known structure to an unknown sequence. Inverse folding is a well-established task in protein research with many recent breakthroughs (Watson et al., 2023). It has recently been adapted for RNA with classical models such as Rosetta (Leman et al., 2020a). Learning based approaches were pioneered by gRNAde (Leman et al., 2020a; Joshi et al., 2024) and other papers specifically address backbone generation (Anand et al., 2024) and protein-binding RNA design (Nori & Jin, 2024).
(Processing): We gather all connected components in our 3D structure database with a size between 15 and 300. Then, we removed identical sequences using a redundancy removal step with CD-HIT with a cutoff of 0.99. We cluster the data on structural similarity with a threshold of 0.9, and split these while ensuring a maximal similarity of 0.5. We also provide datasets and splits from gRNAde (Joshi et al., 2024), that were obtained in a similar manner. Our dataset differs by allowing multi-chain systems, a stricter size cutoff, and a duplicates filter, leading to a reduced dataset size. This task is evaluated with sequence recovery.
4.3 RNA-CM: Dynamic function modulation.
(Context): In addition to the four canonical nucleotides, more than 170 modified nucleotides were found to be integrated in RNA polymers (Boccaletto et al., 2022). Multiple functions of a diverse range of ncRNA have been shown to be directly dependent on such chemical modifications (Roundtree et al., 2017). We propose to detect such chemical modifications from subtle perturbations of the canonical geometry of the RNA backbones. Training a predictor on RNA structures is not the most straightforward way of detecting chemical modifications, but may uncover function-modulating modifications that hitherto went unnoticed.
(Processing): To build this task, we start from the whole dataset, partition it into connected components and apply our size filter. Then, we filter for systems that include modified residues, relying on PDB annotations that flag such modified residues. We apply our default redundancy removal and splitting strategies, which results in 185 data points. Performance is evaluated with accuracy and AuROC.
4.4 RNA-Prot: Biological complex modeling.
@@ 4.2 RNA-IF: Molecular Design
(Context): RNAs and proteins often bind to form a functional complex. Such complexes are involved in crucial cell processes, such as post-transcriptional control of RNAs (Glisovic et al., 2008). RNA structure is often, but not always, heavily disrupted upon interaction with a protein. We expect this task to be an easier version of the RNA-CM task.
(Processing): To build this task, we start from the bgsu non-redundant dataset, partition it into connected components and apply our size and resolution filters. We remove systems originating from ribosomes that are fully encapsulated in a complex. Then, we retain only RNA interacting with a protein and apply our default redundancy removal and splitting strategies, resulting in 1251 data points. Performance is evaluated with accuracy and AuROC.
@@
Further tasks aim to address RNA small-molecule drug-design. RNA is increasingly recognized as a promising family of targets for the development of novel small molecule therapeutics  (Falese et al., 2021; Haga & Phinney, 2023; Disney, 2019; Abulwerdi et al., 2019). Being able to target RNA would drastically increase the size of the druggable space, and propose an alternative for overused protein targets in scenarios where they are insufficient. In addition, they represent a therapeutic avenue in pathologies where protein targets are absent, such as in triple-negative breast cancer (Xu et al., 2020).
4.5 RNA-Site: Drug target detection.
@@
In addition to providing the existing dataset existing in the literature, we propose a larger, up-to-date dataset. We start by following similar steps as for RNA-Prot (without removal of ribosomal systems). We then include a drug-like filter on the small-molecule side. Finally, we include a protein-content filter, removing systems with more than 10 protein atom neighbors, ensuring that the binding is modulated by RNA only. We use the default redundancy removal and splitting, resulting in 458 systems. The predictions are evaluated using accuracy and AuROC.
@@ 4.6 RNA-Ligand: Pocket categorization
(Processing): Starting with similar steps as the RNA-Site, we obtain a set of structures that display RNAs in interaction with one or more drug-like small molecules. We then proceed to extracting the context of the binding pocket by seeding two rounds of breadth-first search with all residues closer than 8Å to an atom of the binder. This results in all binding pockets in our database and we group these pockets with a sequence clustering.
Then, our aim is to find the most frequent ligands binding in non-redundant pockets. To that end, we gather a dataset of binding site clusters that bind to one ligand only. We then compute the most frequent ligands among this dataset and retain only the top three. Subsequently, we discard all binding events to other small molecules, and add clusters that bind only one of the top three ligands to our dataset. This dataset is split on structural similarity and evaluated on MCC and AuROC.
@@ 4.7 RNA-VS: Drug screening
(Processing): The dataset is reproduced from the RNAmigos2 (Carvajal-Patiño et al., 2025) paper. The authors curated a list of binding sites in a similar fashion to the one described in the RNA-Site task and clustered the sites using RMScore (Zheng et al., 2019). All binders found for each cluster are retained as positive examples, and a set of drug-like chemical decoys are added as negative partners. Molecular docking scores are computed with rDock (Ruiz-Carmona et al., 2014) on all binding site-small molecule pairs. The metric used is the AuROC.
@@ 5 Implementation
Next, we briefly showcase the use of our framework for a simple programmatic access to proposed tasks. In Figure 1, we show how practitioners can access our datasets from Python code, automatically downloading them from Zenodo, choosing a representation (in this example a Pytorch Geometric graph) and directly accessing the data in a simple and reproducible fashion. Additionally, by passing a single flag to our loaders, users can choose to execute all processing logic described here to build to build each task from scratch (end-to-end reproducibility).
@@
Figure 1: Obtaining a machine learning-ready split dataset only requires a few lines of code.
@@ 6 Experiments
We hereby explicitly state that we limit the scope of this work to the creation of a useful deep learning library for RNA 3D structure based modeling. We refrain from suggesting novel model architectures; models discussed in this work uniquely serve the goal of illustrating the practical applicability of our library. Within these constraints, we highlight the utility of our benchmark by training a simple relational graph convolutional networks (RGCN) on our proposed tasks. The performance of this baseline is reported in Table 3. Additional details on the experimental setup and results across further metrics can be found in supplementary Section B. These results provide an initial baseline for future model developments on the introduced tasks.
Task Metric Score
RNA_Ligand AuROC 0.6751
RNA_CM Balanced Accuracy 0.6615
RNA_Site Balanced Accuracy 0.6309
RNA_Prot Balanced Accuracy 0.6254
RNA_IF Sequence Recovery 0.3523
RNA_VS AuROC 0.8550
RNA_GO Jaccard 0.3167
Table 3: Selected representative performance metric for each task.
@@
Furthermore, we compare our RGCN approach to existing methods on the TR60/TE18 dataset split (Su et al., 2021) for the RNA-Site task and the dataset and splits from Joshi et al. (2024) for the RNA-VS task. Performance metrics are provided in Supplementary Tables A.2 and A.3. Despite using a relatively simple model, our approach achieves competitive results, aligning with recent methods while falling short of state-of-the-art performance.
@@ 7 Discussion
We have introduced a versatile and modular library designed to advance the application of deep learning to RNA structural analysis. By providing a suite of diverse and well-defined datasets and splits, our library enables robust benchmarking and facilitates the development of innovative computational models. It also ensures reproducibility and promotes standardized evaluation, fostering confidence in computational findings.
In the future, the development of additional tasks, such as the assessment of structural models (Townshend et al., 2021b) and a greater focus on RNA embeddings, presents exciting opportunities.
As the understanding of the dynamic nature of biological macromolecules evolves, so may the preferred representation for RNA. Our tool’s built-in 2.5D-graph representations offer extensive adaptability and have been used to develop recent competitive RNA structure-based models (Wang et al., 2024). Nevertheless, the library can easily be extended to incorporate new or improved representations as they emerge.
@@
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix
Appendix A Details about datasets
In this appendix, we first detail more extensively the datasets and data preprocessing applied (Section A). Then, we provide additional results of the models tested on our tasks together with details about training and hyperparameters used to perform the experiments (Section B). Eventually, we provide the code defining RNA-CM task, which provides a model to introduce new tasks (Section LABEL:sup:sec:code).
A.1 Datasets and general preprocessing
@@
Figure A.1: Distribution of RNAs or RNA chunks sizes
In order to work at a more biologically relevant scale, we partition the raw RNAs from PDB files into connected components in order to use these connected components as samples in our different machine learning tasks (except RNA-Ligand in which the data are partitioned in a different way, into binding pockets rather than into connected components). Figure 1(c) represents the distribution of the number of residues by connected component chunks.
We remove the RNA components of insufficient size for structure-based machine learning, as well as very large components which negatively impact the computational performance of the data loading as well as model training. To this end, we filter RNA chunks and only keep the ones having between 15 and 300 nucleotides. Figure 1(c) displays the distribution of RNA chunks sizes after filtering.
The redundancy removal process we apply is the following. First, we perform sequence-based clustering based on a similarity threshold above which RNA fragments are clustered together (relying on sequence similarity metric CD-HIT (Fu et al., 2012)). Then, within each cluster, we select the sample with the highest resolution. Then we perform structure-based clustering and structure-based redundancy removal (relying on structural similarity metric US-align (Zhang et al., 2022a)) following the same procedure. Afterwards, when instantiating the splitters, we perform structure-based clustering with a different threshold to define the clusters which will be required to be grouped either in train, val or test set (which we name ”splitting clustering”). Since we only select one representative sample per cluster, the number of clusters prefigures the number of samples of the final dataset. Therefore, a tradeoff is to be considered between having a large amount of data and discarding redundant RNAs.
For this reason, we study the impact of both the redundancy removal and the splitting clustering threshold on the number of clusters generated in the case of RNA-CM task. Results are reported in figure A.2. Figure 2(a) displays, for 4 different sequence-based redundancy removal thresholds (each with a different color), the scatter plot of the number of clusters obtained based on the structure-based splitting clustering threshold. Here, the redundancy removal thresholds 0.90, 0.80 and 0.70 give the same plot since CD-HIT similarity values are strongly concentrated around 0.5 and 1. We finally choose 0.90. Figure 2(b) displays, for 4 different structure-based redundancy removal thresholds (each with a different color), the scatter plot of the number of clusters obtained based on the structure-based splitting clustering threshold. In all cases, the structure-based redundancy removal is performed after a sequence-based redundancy removal using a 0.90 threshold. In all the experiments, the splitting clustering is performed based on US-align similarity (structural similarity).
@@
Figure A.2: Number of clusters after similarity clustering based on the splitting clustering threshold
A.2 Task-specific data preprocessing
In this section, we provide additional details about the tasks involving a specific data preparation process.
A.2.1 RNA-Ligand
In this task, we selected the three most frequent ligands which were neither modified RNA or DNA residues nor modified protein residues, following the ligand definition proposed by Hariboss (Panei et al., 2022). We only retained the binding pockets binding to one of these three ligands in order to ensure the amount of binding pockets binding to each ligand would be significant enough to enable learning in a multi-class classification framework. The three ligands retained are paromomycin (called PAR in the PDB nomenclature), gentamycin C1A (LLL) and aminoglycoside TC007 (8UZ). Their structures are displayed in Figure A.3.

### v2

@@ 6. RNA-Ligand Pocket Multi-class (3) 203-43-44
4.4 RNA-Prot: Biological Complex Modeling
@@
(Context): RNAs and proteins often bind to form a functional complex. Such complexes are involved in crucial cell processes, such as post-transcriptional control of RNAs (Glisovic et al., 2008). RNA structure is often, but not always, heavily disrupted upon interaction with a protein.
(Processing): To build this task, we start from the bgsu non-redundant dataset, partition it into connected components and apply our size and resolution filters. We remove systems originating from ribosomes that are fully encapsulated in a complex. Then, we retain only RNA interacting with a protein and apply our default redundancy removal and splitting strategies, yielding 1251 data points.
@@
RNA is increasingly recognized as a promising family of targets for the development of novel small molecule therapeutics  (Falese et al., 2021; Haga and Phinney, 2023; Disney, 2019; Abulwerdi et al., 2019). Being able to target RNA would drastically increase the size of the druggable space, and propose an alternative for overused protein targets in scenarios where they are insufficient. In addition, they represent a therapeutic avenue in pathologies where protein targets are absent, such as in triple-negative breast cancer (Xu et al., 2020). The following tasks aim to address questions in RNA small-molecule drug-design.
4.5 RNA-Site: Drug Target Detection
@@
In addition to providing the existing dataset existing in the literature, we propose a larger, up-to-date dataset. We start by following similar steps as for RNA-Prot (without removal of ribosomal systems). We then include a drug-like filter on the small-molecules. Finally, we include a protein-content filter, removing systems with more than 10 protein atom neighbors, ensuring that the binding is modulated by RNA only. We use the default redundancy removal and splitting, resulting in 458 systems.
@@ 4.6 RNA-Ligand: Pocket Categorization
(Processing): Starting with similar steps as the RNA-Site, we obtain a set of structures that display RNAs in interaction with one or more drug-like small molecules. We then proceed to extract the context of the binding pocket by seeding two rounds of breadth-first search with all residues closer than 8Å to an atom of the binder. This results in all binding pockets in our database and we group these pockets with a sequence clustering.
Then, our aim is to find the most frequent ligands binding in non-redundant pockets. To that end, we gather a dataset of binding site clusters that bind to one ligand only. We then compute the most frequent ligands among this dataset and retain only the top three. Subsequently, we discard all binding events to other small molecules, and add clusters that bind only one of the top three ligands to our dataset and split based on structural similarity.
@@ 4.7 RNA-VS: Drug Screening
(Processing): The dataset is reproduced from the RNAmigos2 (Carvajal-Patiño et al., 2025) paper. The authors curated a list of binding sites in a similar fashion to the one described in the RNA-Site task and clustered the sites using RMScore (Zheng et al., 2019). All binders found for each cluster are retained as positive examples, and a set of drug-like chemical decoys are added as negative partners. Molecular docking scores are computed with rDock (Ruiz-Carmona et al., 2014) on all binding site-small molecule pairs.
@@ 5 Implementation
Next, we briefly showcase the use of our framework for a simple programmatic access to proposed tasks. In Figure 2, we show how practitioners can access our datasets from Python code, automatically downloading them from Zenodo, choosing a representation (in this example a Pytorch Geometric graph) and directly accessing the data in a simple and reproducible fashion. Additionally, by passing a single flag (recompute=True) to the task fetcher, users can choose to execute all processing logic from scratch (end-to-end reproducibility).
@@
Figure 2: Obtaining a machine learning-ready split dataset requires only a few lines of code.
@@ 6 Experiments and Results
We hereby explicitly state that we limit the scope of this work to the creation of a useful deep learning library for RNA 3D structure based modeling. We refrain from suggesting novel model architectures; models discussed in this work uniquely serve the goal of illustrating the practical applicability of our library. However, we provide a baseline performance for each of our proposed tasks as a reference for future works and conduct some investigations highlighting rnaglib’s use in understanding and building RNA structure-function models. All reported results are conducted over three seeds and code is available from github.com/BorgwardtLab/rnaglib-experiments.
6.1 Baseline Performance
The baseline performance for each of our proposed tasks is achieved by training Relational Graph Convolutional Networks (RGCN) on graph representations of the tasks’ underlying datasets. We choose to use 2.5D graphs, whose nodes represent residues and edges represent both canonical and non-canonical interactions. This approach was previously shown to outperform a secondary structure based alternative in a drug-discovery task (Oliver et al., 2020). We report our results in Table 2 and additional details on the experimental setup, including hyperparameters, and further metrics can be found in the supplementary Section A.2. We observe that on all tasks, our machine learning models are able to capture a meaningful signal that outperforms naive approaches, always predicting the majority class, validating the relevance of our tasks for benchmarking purposes.
Table 2: Baseline performance reported using our recommended representative metric for each task. We compare to always predicting the majority class.
Task Metric Majority Score
RNA-GO Jaccard 0.2 0.396 $\pm$ 0.023
RNA-IF Sequence Recovery 0.28 0.372 $\pm$ 0.005
RNA-CM Balanced Accuracy 0.5 0.633 $\pm$ 0.007
RNA-Prot Balanced Accuracy 0.5 0.556 $\pm$ 0.004
RNA-Site Balanced Accuracy 0.5 0.655 $\pm$ 0.012
RNA-Ligand AuROC 0.33 0.650 $\pm$ 0.124
RNA-VS AuROC 0.5 0.759 $\pm$ 0.005
@@ 6.1 Baseline Performance
Furthermore, we compare our RGCN approach to existing methods on the TR60/TE18 dataset split (Su et al., 2021) for the RNA-Site task and the dataset and splits from Joshi et al. (2024) for the RNA-VS task. Performance metrics are provided in Supplementary Tables 4 and 5. Despite using a significantly simpler model, our approach achieves competitive results, aligning with recent methods while falling short of state-of-the-art performance.
6.2 Redundancy and Splitting
To showcase rnaglib’s use in choosing suitable splitting approaches, we compare the three main splitting approaches it provides and present results in Figure 3(a). As expected, training a model on randomly split data leads to inflated results in some cases. However, the difference is not clear cut 3(a). Since our tasks’ datasets already pass a redundancy removal step, we postulate that the reduced redundancy mitigates the effect of leakage. Indeed, forgoing the redundancy removal for an example task (RNA-Site) markedly increases the performance difference between similarity-based and random splitting approaches.
rnaglib permits setting different structure and sequence similarity thresholds for splitting, depending on use case and data availability. In Figure 3(b), we plot the performance of our baseline model on the redundant version of RNA-Site, varying the similarity threshold used. Our results show that more stringent splitting leads to a decreased, more realistic performance. The trade-off between leakage and learning signal can thereby be tweaked by practitioners. We point out that structure and sequence thresholds are not directly comparable, since their calculation differs. Setting a threshold of 1 results in isolated points for our structure similarity metric, resulting in comparable results to random splitting. This is in contrast with our sequence-based similarity metric, whose value is exactly one for highly similar structures. Using either a non-redundant dataset or rigorous splitting, alleviates the risk of data leakage, using both maximizes the reliability of the results.
(a) Figure 3a: Redundancy removal has a dampening effect on data leakage across splitting strategies.
(b) Figure 3b: In a redundant setting, stricter splitting thresholds reduce data leakage, shown for RNA-Site.
6.3 Model Benchmarking
We provide two case studies on selected tasks elucidating how rnaglib can assist in choosing modeling modalities. We investigate the impact of increasing the depth of our model, as well as the effect of using different representations. We focus on RNA-Site, RNA-CM, and RNA-Prot.
First, we assess the impact of varying the number of message passing layers to understand the effect of varying structural context size on model performance. We train models of various depths (number of layers ranging from two to six) on each task of these three tasks, tuning only the learning rate which we adjust for optimal convergence, keeping the remaining hyperparameters as in Supplementary Table 3. Results are presented in Figure 4(a). While using four layers generally results in a good performance, the best value varies from one task to another. For example, predicting protein binding sites (RNA-Prot) might require less structural context than other tasks.
[…]
(a) Figure 4a: RGCN depth
(b) Figure 4b: Structural detail in input RNAs
Next, we assess various types of graph representations for RNA structures. We contrast 2.5D RNA graphs, whose edges include both canonical and non-canonical interactions, rnaglib’s default representation, with 2D RNA graphs (i.e. canonical interactions only). Within the latter, we consider a 2D setting, where all canonical interactions are encoded as one edge type, and a 2D+ setting, where canonical Watson-Crick base-pairing and backbone covalent interactions are distinguished (the backbone’s directionality results in a total of three edge types). We compare these three representations in Figure 4(b) and observe that including non canonical interactions results in a performance increase across all tasks. These results are in accordance with those presented in Oliver et al. (2020).
@@ 7 Discussion
We have introduced a versatile and modular library designed to advance the application of deep learning to RNA structural analysis. By providing a suite of diverse and well-defined datasets and splits, our library enables robust benchmarking and facilitates the development of innovative computational models. It ensures reproducibility and promotes standardized evaluation, fostering confidence in computational findings. Its flexibility allows researchers to easily build additional tasks specific to their research needs in a wide range of RNA structural studies.
In the future, the development of additional tasks, such as the assessment of structural models (Townshend et al., 2021c) and a greater focus on RNA embeddings, presents exciting opportunities. As the understanding of the dynamic nature of biological macromolecules evolves, so may the preferred representation for RNA. Our tool’s built-in 2.5D-graph representations offer extensive adaptability and have been used to develop recent competitive RNA structure-based models (Wang et al., 2024). Nevertheless, the library can easily be extended to incorporate new or improved representations as they emerge.
@@
Appendix A Technical Appendices and Supplementary Material
A.1 Details about datasets
In this appendix, we first detail more extensively the datasets and data preprocessing applied (Section A.1). Then, we provide additional results of the models tested on our tasks together with details about training and hyperparameters used to perform the experiments (Section A.2).
A.1.1 Datasets and general preprocessing
@@
Figure 5: Distribution of RNAs or RNA chunks sizes
A single system, such as rRNA, can gather several independent, non-interacting RNA chains in a single PDB file. In order to work at a more biologically relevant scale, we partition the raw RNAs from PDB files into connected components. Figure 5(c) represents the distribution of the number of residues by connected components.
We remove the RNA components of insufficient size for structure-based machine learning, as well as very large components which negatively impact the computational performance of the data loading as well as model training. To this end, we filter RNA chunks and only keep the ones having between 15 and 500 nucleotides. Figure 5(c) displays the distribution of RNA chunks sizes after filtering.
The redundancy removal process we apply is the following. First, we perform sequence-based clustering based on a similarity threshold above which RNA fragments are clustered together (relying on a sequence similarity metric computed using CD-HIT [Fu et al., 2012] clustering program). Then, within each cluster, we select the sample with the highest resolution. Then we perform structure-based clustering and structure-based redundancy removal (relying on the TM-Score, a structural similarity metric, computed using US-align platform [Zhang et al., 2022a]) following the same procedure. Afterwards, when instantiating the splitters, we perform structure-based clustering with a different threshold to define the clusters which will be required to be grouped either in train, val or test set (which we call "splitting clustering"). Since we only select one representative sample per cluster, the number of clusters prefigures the number of samples of the final dataset. Therefore, a tradeoff is to be considered between having a large amount of data and discarding redundant RNAs.
For this reason, we study the impact of both the redundancy removal and the splitting clustering threshold on the number of clusters generated in the case of RNA-CM task. Results are reported in figure 6. Figure 6(a) displays, for 4 different sequence-based redundancy removal thresholds (each with a different color), the scatter plot of the number of clusters obtained based on the structure-based splitting clustering threshold. Here, the redundancy removal thresholds 0.90, 0.80 and 0.70 give the same plot since CD-HIT similarity values are strongly concentrated around 0.5 and 1. The chosen value is 0.90. Figure 6(b) displays, for 4 different structure-based redundancy removal thresholds (each with a different color), the scatter plot of the number of clusters obtained based on the structure-based splitting clustering threshold. In all cases, the structure-based redundancy removal is performed after a sequence-based redundancy removal using a 0.90 threshold. In all the experiments, the splitting clustering is performed based on US-align similarity (structural similarity).
@@
Figure 6: Number of clusters after similarity clustering based on the splitting clustering threshold
A.1.2 Task-specific data preprocessing
We provide additional details about the tasks involving a specific data preparation process.
RNA-Ligand
We selected the three most frequent ligands which were neither modified RNA or DNA residues nor modified protein residues, following the ligand definition proposed by Hariboss [Panei et al., 2022]. We only retained the binding pockets binding to one of these three ligands in order to ensure the amount of binding pockets binding to each ligand would be significant enough to enable learning in a multi-class classification framework. The three ligands retained are paromomycin (called PAR in the PDB nomenclature), gentamicin C1A (LLL) and aminoglycoside TC007 (8UZ). Their structures are displayed in Figure 7.